# Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v1

@@ A.9.1 How does this affect our results?
Within our experiments we normalise the training and testing data, as well as sample random observations from the dataset. Because of this, observations predicted during in-context learning will be likely not have been memorised, since they are normalised based on a random training data subset, and are selected from random rows of the dataset. As shown, although the language model is able to complete the first few rows of the datasets with some success, it is unable to complete random rows (Table 4). Therefore, we hypothesise that any memorisation of the tested datasets has little impact on the in-context predictions.
@@ A.10 Task descriptions with more information
To understand how the task description affects the performance of the priors elicited from the language model, we studied the synthetic task, $y=2x_{1}-x_{2}+x_{3}$, and described the relationship between features and targets in varied detail. We additionally test an uninformative prior to contextualise the results. For each of the levels of detail shown in Figure 13, we included the following text in the task description that was given to the language model:
@@
What is unexpected, however, is that for this task, using an uninformative prior ($\theta\sim\mathcal{N}(\theta|0,1)$) provides better posterior performance than specifying to the language model that the relationship between the target and features is linear (Figure 13). This is explained by inspecting the elicited prior distribution when being given this information, which shows that the language model responded with a similar mean and standard deviation for a Gaussian for each of the $100$ variations of prompts:
@@
Interestingly, although “Three feature relationships” and “Full equation” contain the same information written differently, the language model provided better prior distributions when given the full equation. Fully specifying the equation is not realistic in a real-world setting where we are not sure about the exact relationship between features and targets, but this provides an example of how we can make the task description as information-dense as possible.
@@
An expert interacting with the language model might be able to provide further information about a task, that improves the elicited priors. In Figure 14 we show the parameter distributions elicited from the language model for the different descriptions we provide. This figure shows that as we provide additional information about each feature, the language model responds by providing a prior distribution that reflects its updated knowledge of the task. This demonstrates the potential of using an expert to construct more detailed task descriptions, which the language model can translate into parameter distributions and supplements the results in Section 3.2.
@@ A.11 How many ways should we describe a task?
Figure 15 shows the energy statistic (Appendix A.4) of samples of the prior distribution elicited using fewer task descriptions compared to samples from the prior distribution elicited using $100$ task descriptions (from the same language model). In particular, we compare a subset of the task descriptions to the full $100$ task descriptions.
We see that for most tasks, a large number of task descriptions is required, since reducing the number of descriptions from $100$ to $49$ produced prior distributions that are significantly different for all datasets tested. Using $81$ task descriptions produces prior distributions that are more similar but still show differences for some datasets (for example $y=2x_{1}-x_{2}+x_{3}$ and Breast Cancer).
It is interesting to see that Breast Cancer benefitted most from increasing the number of task descriptions, since this dataset contains the largest number of features. It suggests that more task descriptions are required for those datasets with high-dimensional data. This makes intuitive sense, since we would expect to require a larger mixture of Gaussian components to better describe the prior over for a higher dimensional dataset.
@@ A.12 Cost of elicitation
The priors elicited from GPT-3.5-turbo incur a cost, as we use OpenAI’s API. To elicit all of the priors used in the experiments in Section 3.1, we were charged \$$\$0.41$ (on average \$$\$0.068$ per dataset). However, we did not make use of the batch API and instead made each API call sequentially. All of the experiments in this work could be implemented using the batch API since no new API call depends on a previous API result, which may reduce cost significantly  [^4] https://openai.com/api/pricing/.
@@ A.13.1 Other OpenAI models
In Figure 16 we also present the posterior results of priors elicited from both GPT-4-turbo and GPT-4o-mini. GPT-4-turbo incurs a significantly higher cost than the GPT-3.5-turbo model used for the main results in this paper, whilst GPT-4o-mini provides a less expensive option.
@@
Figure 16 demonstrates that in most cases, GPT-4-turbo provides priors that are significantly better than GPT-3.5-turbo but falls behind in the regression tasks. On Breast Cancer, UTI, and Wine Quality, the priors elicited from GPT-4-turbo allow for a significantly better posterior accuracy, and unlike GPT-3.5-turbo, provide an informative prior for Heart Disease. However, for both the regression tasks (the synthetic task and the California Housing task), the GPT-3.5-turbo priors allow for a better mean squared error (MSE) than GPT-4-turbo.
@@ A.13.2 Open source language models
In contrast to the OpenAI models in Appendix A.13.1, these models are publicly available and open-source after being granted access. This means they can be used without the need for connecting to an online API, and when locally hosted, all data and inference can be performed on device, enabling the use of our framework within a closed and secure environment.
The experiments presented in this section use the same task descriptions as given to GPT-3.5-turbo, except that because the Llama models often fail to return a JSON structure, we suffix the prompt with an additional instruction to return a JSON. When sampling the posterior with the elicited priors, we use the same methods and options as given in A.5.2.
Figure 17 illustrates the inconsistent posterior accuracy of the priors elicited from the Llama family of language models tested. These models are significantly smaller in capacity to GPT-3.5-turbo, justifying their inconsistent elicited priors, but in some of the datasets, they still outperform the uninformative prior.
Specifically for Breast Cancer, the 3B parameter Llama model provides an intermediate accuracy between an uninformative prior and those elicited from GPT-3.5-turbo, and for the synthetic task $y=2x_{1}-x_{2}+x_{3}$ and California Housing, the 8B parameter Llama provides priors of equal quality to GPT-3.5-turbo. On the other hand, for the UTI, Wine Quality, and Heart Disease datasets, the Llama models both provide prior distributions that are of worse quality than the uninformative prior. It is possibly not surprising that the Llama-3.2-3B-Instruct provided priors that were equal or worse than Llama-3.1-8B-Instruct in all cases except for predicting Breast Cancer since these models differ significantly in size.
@@
Qwen-2.5-14B-Instruct is similarly inconsistent in the posterior performance that the priors provide. For Heart disease, it achieves a greater posterior accuracy than all models tested, except for the 70B Llama model, Llama-3.1-70B-Instruct-BF16. For the UTI task, after $10$ demonstrations are seen, it performs comparably to Llama-3.1-70B-Instruct-BF16, and better than GPT-3.5-turbo. On Wine Quality it achieves the greatest posterior accuracy after $20$ observed data points. However, on the California Housing, Breast Cancer, and the synthetic task, it performs worse than GPT-3.5-turbo, but is likely smaller in capacity. Overall, Qwen-2.5-14B-Instruct does not provide consistent performance across the tasks, but has the benefit of performing similarly to Llama-3.1-70B-Instruct-BF16 on many tasks with a significantly smaller compute requirement.

### v2

@@ A.9.1 How does this affect our results?
Within our experiments, we normalise the training and testing data, as well as sample random observations from the dataset. Because of this, observations predicted during in-context learning will likely not have been memorised, since they are normalised based on a training data subset and are selected from random rows of the dataset. As shown, although the language model can complete the first few rows of the datasets with some success, it is unable to complete random rows (Table 4). Therefore, we hypothesise that any memorisation of the tested datasets has little impact on the in-context predictions.
@@ A.10 Task descriptions with more information
To understand how the task description affects the performance of the priors elicited from the language model, we studied the synthetic task, $y=2x_{1}-x_{2}+x_{3}$, and described the relationship between features and targets in varied detail. We additionally test an uninformative prior to contextualise the results. For each of the levels of detail shown in Figure 13, we included the following text in the task description that was provided to the language model:
@@
However, what is unexpected is that for this task, using an uninformative prior ($\theta\sim\mathcal{N}(\theta|0,1)$) provides better posterior performance than specifying to the language model that the relationship between the target and features is linear (Figure 13). This is explained by inspecting the prior distribution elicited when given this information, which shows that the language model responded with a similar mean and standard deviation for a Gaussian for each of the $100$ variations of the prompts:
@@
Interestingly, although “Three feature relationships” and “Full equation” contain the same information written differently, the language model provided better prior distributions when given the full equation. Fully specifying the equation is not realistic in a real-world setting where we are not sure about the exact relationship between features and targets, but this provides an example of how we can make the task description as information dense as possible.
@@
An expert interacting with the language model might be able to provide further information about a task that improves the elicited priors. In Figure 14 we show the parameter distributions elicited from the language model for the different descriptions we provide. This figure shows that, as we provide additional information about each feature, the language model responds by providing a prior distribution that reflects its updated knowledge of the task. This demonstrates the potential of using an expert to construct more detailed task descriptions, which the language model can translate into parameter distributions and supplements the results in Section 3.2.
@@ A.11 How many ways should we describe a task?
Figure 15 shows the energy statistic (Appendix A.4) of samples from the prior distribution elicited using fewer task descriptions compared to samples from the prior distribution elicited using $100$ task descriptions (from the same language model). In particular, we compare a subset of task descriptions with the full $100$ task descriptions.
We see that for most tasks, a large number of task descriptions is required, since reducing the number of descriptions from $100$ to $49$ produced prior distributions that are significantly different for all datasets tested. Using $81$ task descriptions produces prior distributions that are more similar but still show differences for some datasets (for example, $y=2x_{1}-x_{2}+x_{3}$ and Breast Cancer).
It is interesting to see that Breast Cancer benefitted most from increasing the number of task descriptions, since this dataset contains the largest number of features. It suggests that more task descriptions are required for those datasets with high-dimensional data. This makes intuitive sense, since we would expect to require a larger mixture of Gaussian components to better describe the prior over for a higher-dimensional dataset.
@@ A.12 Cost of elicitation
The priors elicited from GPT-3.5-turbo incur a cost, as we use OpenAI’s API. To elicit all of the priors used in the experiments in Section 3.1, we were charged \$$\$0.41$ (on average \$$\$0.068$ per dataset). However, we did not make use of the batch API and instead made each API call sequentially. All experiments in this work could be implemented using the batch API since no new API call depends on a previous API result, which may significantly reduce costs  [^4] https://openai.com/api/pricing/.
@@ A.13.1 Other OpenAI models
In Figure 16 we also present the posterior results of the priors elicited from both GPT-4-turbo and GPT-4o-mini. GPT-4-turbo incurs a significantly higher cost than the GPT-3.5-turbo model used for the main results in this paper, whilst GPT-4o-mini provides a less expensive option.
@@
Figure 16 demonstrates that in most cases, GPT-4-turbo provides priors that are significantly better than GPT-3.5-turbo but falls behind in the regression tasks. On Breast Cancer, UTI, and Wine Quality, the priors elicited from GPT-4-turbo allow for a significantly better posterior accuracy, and unlike GPT-3.5-turbo, provide an informative prior for Heart Disease. However, for both regression tasks (the synthetic task and the California Housing task), the GPT-3.5-turbo priors allow for a better mean squared error (MSE) than GPT-4-turbo.
@@ A.13.2 Open source language models
In contrast to the OpenAI models in Appendix A.13.1, these models are publicly available and open source after being granted access. This means that they can be used without the need for connecting to an online API, and when locally hosted, all data and inference can be performed on device, enabling the use of our framework within a closed and secure environment.
The experiments presented in this section use the same task descriptions as given to GPT-3.5-turbo, except that because Llama models often fail to return a JSON structure, we suffix the prompt with an additional instruction to return a JSON. When sampling the posterior with the elicited priors, we use the same methods and options as given in A.5.2.
Figure 17 illustrates the inconsistent posterior accuracy of the priors elicited from the Llama family of tested language models. These models are significantly smaller in capacity to GPT-3.5-turbo, justifying their inconsistent elicited priors, but in some of the datasets, they still outperform the uninformative prior.
Specifically for Breast Cancer, the 3B parameter Llama model provides an intermediate accuracy between an uninformative prior and those elicited from GPT-3.5-turbo, and for the synthetic task $y=2x_{1}-x_{2}+x_{3}$ and California Housing, the 8B parameter Llama provides priors of equal quality to GPT-3.5-turbo. However, for the UTI, Wine Quality, and Heart Disease datasets, both Llama models provide prior distributions that are worse quality than the uninformative prior. It is possibly not surprising that the Llama-3.2-3B-Instruct provided priors that were equal or worse than Llama-3.1-8B-Instruct in all cases except for predicting Breast Cancer since these models are significantly smaller in size.
@@
Qwen-2.5-14B-Instruct is similarly inconsistent in the posterior performance that the priors provide. For Heart Disease, it achieves a greater posterior accuracy than all models tested, except for the 70B Llama model, Llama-3.1-70B-Instruct-BF16. For the UTI task, after seeing $10$ demonstrations, it performs comparably to Llama-3.1-70B-Instruct-BF16, and better than GPT-3.5-turbo. On Wine Quality it achieves the greatest posterior accuracy after $20$ observed data points. However, on the California Housing, Breast Cancer, and the synthetic task, it performs worse than GPT-3.5-turbo, but is likely smaller in capacity. Overall, Qwen-2.5-14B-Instruct does not provide consistent performance across tasks, but has the benefit of performing similarly to Llama-3.1-70B-Instruct-BF16 on many tasks with a significantly smaller compute requirement.